OpenStudy (anonymous):

how to simplify the third root of x, to the twelfth power I know it equals x^4 but I want to know HOW to get the answer.

Nnesha (nnesha):

\[\huge\rm ( \sqrt[3]{x})^{12}\] like this right

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so you can change root to exponent \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\]

Nnesha (nnesha):

so \[\huge\rm \sqrt[3]{x} = x^{?}\ ] how would you change 3rd root to exponent ?

OpenStudy (anonymous):

I didn't get the visual

Nnesha (nnesha):

hmm latex is not working  how would you change 3rd too to exponent ?